OpenStudy (anonymous):

The slope of any line perpendicular to , where A(0, -3) and B(4, 9) are two points on the line, is m =

OpenStudy (anonymous):

first lets find the slope slope = (y2 - y1) / (x2 - x1) (0,-3) x1 = 0 and y1 = -3 (4,9) x2 = 4 and y2 = 9 now lets sub slope = (9 - (-3) / (4 - 0) slope = (9 + 3) / 4 slope = 12/4 which reduces to 3 so the slope given is 3, however, we want the slope for a perpendicular line. All that means is take the slope given, and " flip " it, then change the sign. This is called the negative reciprocal. So the negative reciprocal of 3 (or 3/1) is -1/3 -- see how I flipped the slope and changed the sign. For a perpendicular line, m has to equal -1/3. any questions ?